Suppose you wish to test if a number cube (die) is loaded or not. If the die is not loaded, the theoretical probabilities for each roll should be:

1

2

3

4

5

6

16 2/3 %    16 2/3 % 16 2/3 % 16 2/3 % 16 2/3 % 16 2/3 %


You roll the die 93 times and come up with the following distribution:

1

2

3

4

5

6

15    13 21 14 15 15


What type of test should be used in this situation and what is the test statistic?

a) χ2 Goodness of Fit; χ2 = 2.548

b) One proportion z test; z = 0.161

c) One proportion z test; z = 1.230

d) χ2 Goodness of Fit; χ2 = 4.074

e) Two proportion z test; z = 0.161

f) None of the above
Answer #1

If a die is rolled each side is equally probable. Probabilty of each side 16.67 % or 1/6 .

Expected count for each side is = 93 * (1/6) = 15.5.

Number on face Observe count Expected count
1 15 15.5
2 13 15.5
3 21 15.5
4 14 15.5
5 15 15.5
6 15 15.5

I shall use 2 goodness of fit.

2 = ( observed - expected ) ^2 / expected = (15- 15.5 )^2 / 15.5 + ... + (15- 15.5 )^2 / 15.5 = 2.548

So, option a) is correct.

Feel free to comment if you need any clarification.

If you find my answer useful then please support me by putting thumbs-up . Thank you.